OpenStudy (anonymous):

Chip planted flowers in a rectangular garden and in a square garden. Which statement about the gardens is true? A. The area of the rectangular garden is greater than the area of the square garden. B. The area of both gardens is the same. C. The perimeter of the square garden is greater than the perimeter of the rectangular garden. D. The perimeter of both gardens is the same.

jhonyy9 (jhonyy9):

for perimeter what is greater 2*(96+70) = ? or 4*80 = ? ??? for area what is greater ? 96*70 = ? or 80*80 = ? ???

OpenStudy (anonymous):

=6400